OpenStudy (anonymous):

In order to go to his friends house, Luis must drive 24 miles south then 10 miles east. how many miles could he save if he could drive the straight-line distance between the 2 houses?

OpenStudy (anonymous):

10^2 + 24^2 = C^2 C = 26 24 + 10 - 26 = 8

OpenStudy (anonymous):

the width of a rectangle is 5 feet and the diagonal is 9 feet. what is the area of the ractangle?

OpenStudy (anonymous):

is it the same concept?

OpenStudy (radar):

Kinda. You solve for the length then calculate area.

OpenStudy (anonymous):

well i did the 5*5 and 9*9 and got 106

OpenStudy (anonymous):

got a whole bunch of decimals as the answer. tthe answers highlighted are 2square root 14 10 square root 14 2 square feet 10 square feet or Not shown

OpenStudy (anonymous):

I guess u cant take the square root of 106

OpenStudy (radar):

\[l=\sqrt{81-25}=\sqrt{56}\] Area A=(5)(7.483)=37.41657387 sq ft

OpenStudy (anonymous):

see that answer isnt on there

OpenStudy (radar):

10 times square root of 14

OpenStudy (radar):

\[\sqrt{56}=\sqrt{14}\sqrt{4}=2\sqrt{14}\]Multipying by 5 gives result ur looking for.

OpenStudy (radar):

new problem Luis Rivera

OpenStudy (radar):

a.

OpenStudy (radar):

\[\sqrt{500}=10\sqrt{5}\]